Jan's account Ballenger is $3800. Of this, $750 is for rent. Jan wants to keep a ballence of at least $500. Write and solve an inequality describing how much she can withdrawal and still leave enough for rent and a $500 ballence.

Answer by stanbon(75887)

Jan's account Ballenger is $3800. Of this, $750 is for rent. Jan wants to keep a balance of at least $500. Write and solve an inequality describing how much she can withdrawal and still leave enough for rent and a $500 balance.
750+500 = 3800 - x
----
x = 3800-(1250)
x = $2550 (amt. she can withdraw)
-------
Cheers,
Stan H.
-----------